If you have completely polarized light of intensity 140 W/m2, what will its intensity be after passing through a polarizing filter with its axis at an 88.5° angle to the light's polarization direction?
Answer #1

Malu's law,

The intensity is proportional to the square of the cosine of the angle.

I new = I (cos theta )^2 = 145 * (cos 88.5)^2 = 0.0959 W/m^2